Three consecutive integers are as such they are taken in increasing order and multiplied by 2, 3, and 4, respectively, they add up to 56. Find these numbers.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us say the three consecutive numbers be x, x+1 and x+2.

Now as per the given question;

2×(x)+3×(x+1)+4×(x+2)=56
9x + 11 = 56
9x = 56-11
9x = 45
x = 45/9
x= 5

Thus, the three consecutive numbers are 5, 6 and 7.

Thanks